# Time-Crystalline Phase in a Single-Band Holographic Superconductor

### v1

@@ V Time-crystalline phase from bulk numerics
1. Solve coupled ODEs (4) and (5) for $\{\psi_{0},\phi_{0}\}$ with boundary conditions (6), given a constant chemical potential. To achieve this, we employ a shooting method from the horizon to the boundary. For stability, we tune the initial condensate value to match the source-free boundary condition $\psi^{(1)}=0$, achieving convergence to within $10^{-4}$ relative error.
@@
As shown in FIG. $1$, we found that for small driving current, the Higgs mode oscillates at twice the driving frequency (harmonic response). For large driving, nonlinearity leads to envelope growth and distortion. In Fig. 2, a finite number of resonant peaks are observed in the spectra of both the 2:1 and sum channels. These subharmonic peaks, in particular, signify the presence of a time-crystalline phase Greilich:2024 .
@@ VI Discussion and Conclusion
Building on this controlled reduction, we have presented a self-contained derivation of boundary effective ordinary differential equations governing driven, homogeneous dynamics of a single-band holographic superconductor, with all coefficients fixed from first principles and appropriate rescalings. A multiple-scale analysis of these equations predicts robust subharmonic (time-crystalline) responses, in quantitative agreement with numerical simulations. The formalism is readily generalizable to multi-band holographic superconductors, as well as to inhomogeneous or finite-size drive geometries. Extensions incorporating momentum-dependent couplings, finite-size effects, or more realistic lattice structures can be implemented within the same framework Wen:2013ufa ; Zhang:2025hkb .
Beyond the minimal setup considered here, several physically motivated directions merit further investigation. Finite-temperature extensions could clarify crossover behavior and the fate of time-crystalline responses near criticality, while coupling to competing orders—such as charge density waves or nematic phases—may generate richer dynamical phase diagrams. Such studies could establish connections to experimentally observed nonequilibrium phenomena in bilayer graphene, moiré materials, and related strongly correlated systems Zhao:2023 . These extensions may also shed light on the interplay between driven superconductivity and other collective modes in complex quantum materials.
More broadly, by applying holographic methods to time-crystalline behavior previously observed in simulated lattice high-$T_{c}$ superconductors, this work highlights the utility of gauge–gravity duality in capturing intrinsically nonperturbative dynamics inaccessible to weak-coupling approaches. The predicted subharmonic responses closely resemble experimental signatures observed in optically pumped cuprates, suggesting that holographic models can serve as effective simulators for ultrafast spectroscopy. Future directions include incorporating quantum noise through stochastic holography to assess stability against decoherence, as well as explicitly modeling coupling to external baths to realize open-system time crystals. Ultimately, these insights may inform the design of quantum devices that exploit broken time-translation symmetry to enhance coherence times or enable novel Floquet engineering protocols.

### v2

@@ V Time-crystalline phase from bulk numerics
Figure 1: (Left) Contour plot of given boundary conditions for fixed $\mu$ and $\psi^{(1)}=0$ in the shooting parameters space. Solutions exist at intersecting points. For higher chemical potential (for instant, $\mu=10$), there exists multiple solutions for excited states. (Right) In our simulation, we use the ground state wavefunctions $\psi(z)$ and $\phi(z)$ for $\mu=2.5$.
@@
1. Solve coupled ODEs (4) and (5) for $\{\psi_{0},\phi_{0}\}$ with boundary conditions (6), given a constant chemical potential. To achieve this, we employ a shooting method from the horizon to the boundary. For stability, we tune the initial condensate value to match the source-free boundary condition $\psi^{(1)}=0$, achieving convergence to within $10^{-4}$ relative error, see Fig. 1.
@@
As shown in FIG. 2, we found that for small driving current, the Higgs mode oscillates at twice the driving frequency (harmonic response). For large driving, nonlinearity leads to envelope growth and distortion. In Fig. 3, a finite number of resonant peaks are observed in the spectra of both the 2:1 and sum channels. These subharmonic peaks, in particular, signify the presence of a time-crystalline phase Greilich:2024 .
@@ VI Discussion and Conclusion
Building on this controlled reduction, we have presented a self-contained derivation of boundary effective ordinary differential equations governing driven, homogeneous dynamics of a single-band holographic superi. conductor, with all coefficients fixed from first principles and appropriate rescalings. A multiple-scale analysis of these equations predicts robust subharmonic (time-crystalline) responses, in quantitative agreement with numerical simulations. The formalism is readily generalizable to multi-band holographic superconductors, as well as to inhomogeneous or finite-size drive geometries. Extensions incorporating momentum-dependent couplings, finite-size effects, or more realistic lattice structures can be implemented within the same framework Wen:2013ufa ; Zhang:2025hkb .
Beyond the minimal setup considered here, several physically motivated directions merit further investigation. Finite-temperature extensions could clarify crossover behavior and the fate of time-crystalline responses near criticality [^3] We thank Jeong:2023las for pointing out their holographic approach on Anderson-Higgs mechanism, where a crossover of Higgs mode was observed at about half $T_{c}$. We do not expect to observe this crossover since we have chosen the fixing chemical potential such that boundary temperature is not far away from $T_{c}$., while coupling to competing orders—such as charge density waves or nematic phases—may generate richer dynamical phase diagrams. Such studies could establish connections to experimentally observed nonequilibrium phenomena in bilayer graphene, moiré materials, and related strongly correlated systems Zhao:2023 . These extensions may also shed light on the interplay between driven superconductivity and other collective modes in complex quantum materials.
More broadly, by applying holographic methods to time-crystalline behavior previously observed in simulated lattice high-$T_{c}$ superconductors, this work highlights the utility of gauge/gravity duality in capturing intrinsically nonperturbative dynamics inaccessible to weak-coupling approaches. The predicted subharmonic responses closely resemble the experimental signatures observed in optically pumped cuprates, suggesting that holographic models can serve as effective simulators for ultrafast spectroscopy. Future directions include incorporating quantum noise through stochastic holography to assess stability against decoherence, as well as explicitly modeling coupling to external baths to realize open-system time crystals. Ultimately, these insights may inform the design of quantum devices that exploit broken time-translation symmetry to enhance coherence times or enable novel Floquet engineering protocols.